OpenStudy (anonymous):

how do i factor 4x^2+9?

OpenStudy (anonymous):

4x² - 9 = (2x - 3)(2x + 3)

zepdrix (zepdrix):

April have you been working with complex (imaginary) numbers in your class?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

and I need the steps of how to get (2x-3)(2x+3)

zepdrix (zepdrix):

No? Hmm, did you put a plus when you meant to put a minus? As Haiqa posted.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then its

OpenStudy (anonymous):

(2x + 3) ^2

zepdrix (zepdrix):

\[\large 4x^2+9\]Does NOT factor into \(\large (2x-3)(2x+3)\) because of the addition betwee the terms. :\

zepdrix (zepdrix):

It factors into \(\large (2x-3i)(2x+3i)\) You sure it wasn't suppose to be subtraction? D: lol

OpenStudy (anonymous):

positive lol

zepdrix (zepdrix):

Ok well anyway, here is a step that will help. We want to rewrite each term as a square. \[\large 4x^2+9 \qquad = \qquad 2^2x^2+3^2 \qquad = \qquad (2x)^2+3^2\]Understand what I did so far?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Since we have the `sum of squares` we can rewrite them as the product of complex conjugates using this rule,\[\large a^2+b^2=(a+bi)(a-bi)\]

OpenStudy (anonymous):

l

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

If you haven't been introduced to imaginary numbers yet, then your book or teacher might actually want you to put "no solution" or something for the answer. But this is the method we want to use if we're understanding the concept :)

OpenStudy (anonymous):

oh wait yeah we did learn about imaginary numbers,

zepdrix (zepdrix):

lol :3

OpenStudy (anonymous):

(2x-3i)(2x+3i)

zepdrix (zepdrix):

Understand how that rule will help us? In our case we have \(\large a=(2x)\) and \(\large b=3\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so would my final answer be (2x-3i)(2x+3i)?

zepdrix (zepdrix):

yes :)

OpenStudy (jhannybean):

do you know WHY it's i though?..

OpenStudy (anonymous):

Thanks so much!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yeah imaginary number